### v1

# A Necessary Step toward Faithfulness: Measuring and Improving Consistency in Free-Text Explanations

Abstract
Faithful free-text explanations are important to ensure transparency in high-stakes AI decision-making contexts, but they are challenging to generate by language models and assess by humans. In this paper, we present a measure for Prediction-EXplanation (PEX) consistency, by extending the concept of weight of evidence. This measure quantifies how much a free-text explanation supports or opposes a prediction, serving as an important aspect of explanation faithfulness. Our analysis reveals that more than 62% explanations generated by large language models lack this consistency. We show that applying direct preference optimization improves the consistency of generated explanations across three model families, with improvement ranging from 43.1% to 292.3%. Furthermore, we demonstrate that optimizing this consistency measure can improve explanation faithfulness by up to 9.7%.
A Necessary Step toward Faithfulness:
Measuring and Improving Consistency in Free-Text Explanations
Lingjun Zhao University of Maryland College Park, Maryland, USA lzhao123@umd.edu                        Hal Daumé III University of Maryland College Park, Maryland, USA hal3@umd.edu
1 Introduction
Explainable AI systems are those that can specify the relationship between an output or prediction and the deductive or nomological process that led to that prediction through logically consistent and empirically grounded means Woodward and Ross (2003).  [^1] Other forms of explanations exist, such as those that aim to provide useful information for better decision making Kayser et al. (2024); Han et al. (2023) or debugging Ribeiro et al. (2016); Lundberg and Lee (2017). Such explanations are often called “faithful” in the sense that they reflect a model’s true reasoning process Jacovi and Goldberg (2020); Lyu et al. (2024), and faithfulness is often seen as crucial for trustworthiness, transparency, and accountability.
Faithfulness, however, is a difficult construct to measure for complex models (such as deep neural networks) precisely because we do not know a priori exactly how a models combines its input features to make a prediction, beyond trivially writing out the entire computation—which is then impenetrable to a person. When an explanation is given in the form of natural language, increasingly common especially in the context of large language models Camburu et al. (2018); Wiegreffe et al. (2022) the challenge increases because humans often conflate faithfulness and plausibility: how convincing it appears Jacovi and Goldberg (2020); Lyu et al. (2023); Wiegreffe et al. (2022). This leads to models that produce plausible yet unfaithful accounts of their reasoning processes Ye and Durrett (2022); Lanham et al. (2023), which risks eroding the very trustworthiness that they aim to improve.
Figure 1: Example of GPT-4 model generating explanations for truthful or deceptive prediction about a hotel review’s authenticity. While each explanation appears plausible on its own, they are inconsistent: the explanation “use a lot of positive language” is used to support both truthful and deceptive predictions.
In this paper, we aim to improve model faithfulness by optimizing the consistency of an explanation with its explanandum (e.g., prediction). If an explanation is to be a faithful presentation of a deductive of nomological process, then it must hold that an explanation $e$ cannot simultaneously support a prediction $y$ and its negation $\lnot y$. For example, in Figure 1, we see that a model uses virtually the same explanation (“use a lot of positive language’) to argue in favor of a hotel review being both authentic and deceptive in this opinion spam classification task.
Our work proceeds in three steps. First, we ask how consistent large language model (LLM)-generated explanations are in practice. To do this, we introduce a measure of consistency, PEX (for Prediction-EXplanation consistency), which can be applied to models that provide probabilities over their generated outputs. PEX leverages the weight of evidence framework Melis et al. (2021); Wod (1985) to quantify how much an explanation speaks in favor of (vs against) a prediction (§ 3.1, § 3.2). On two datasets—TripAdvisor hotel review and Amazon product review spam detection Ott et al. (2013); Hussain et al. (2020)—We find that 62%–86% of explanations from Llama-2 Touvron et al. (2023), Mistral Mistral (2023), and Yi-1.5 Young et al. (2024) are inconsistent, logically implying that they are also unfaithful (§ 6.1).
Second, we ask whether we can use our consistency measure PEX to train a model to produce more consistent—and therefore less unfaithful—explanations. We show how to apply both supervised fine-tuning and direct preference optimization (DPO) Rafailov et al. (2024) to refine pretrained LLMs (§ 3.4). For DPO, we sample explanations from LLMs and rank them according to PEX; those that score highly are treated as “preferred” in the optimization, and those that score lowly are treated as “dispreferred.” Experimentally, we show this DPO-based approach significantly improves on supervised fine-tuning and can improve explanation consistency by 43.1%-292.3% (§ 6.2).
It is, of course, possible that the fine-tuning leads to more consistent explanations, but not more faithful explanations. And so finally we measure whether the explanations that we optimized for consistency lead to improved faithfulness. We construct a proxy measure of faithfulness by assuming that a more faithful explanation should be more useful for an external observer to predict a model’s behavior (Lyu et al., 2024, see Section 3.4.2). Using this idea, we adopt a simulatability-based explanation faithfulness evaluation method Pruthi et al. (2022) (§ 4) and show that optimizing for PEX improves faithfulness by 1.5% to 9.7% (§ 6.3).
2 Related Work
Challenges in ensuring faithfulness of free-text explanations.
Different from interpretable AI Wallace et al. (2018); Selvaraju et al. (2017); Ribeiro et al. (2016); Briakou et al. (2023), generating explanations is a procedure to explicitly explaining model decisions to people Miller (2019). A faithful explanation should accurately reflect the reasoning process behind the model’s prediction Jacovi and Goldberg (2020); Lyu et al. (2024); Ribeiro et al. (2016). For models that first make a prediction with a standard black-box predictor and then justify the prediction with an explainer Camburu et al. (2018); Park et al. (2018); Wu and Mooney (2019), there is no guarantee for the explanation faithfulness. For models that first generate an explanation and then provided as the only input to the predictor, the explanations can still be self-inconsistent Camburu et al. (2020); Zhou et al. (2023), or optimized in terms of plausibility instead of faithfulness Kumar and Talukdar (2020). Most methods that jointly explain and make prediction Rajani et al. (2019); Narang et al. (2020); Ling et al. (2017); Jung et al. (2022); Ramnath et al. (2024) also do not ensure faithfulness, as models may ignore explanation during prediction. For chain-of-thought style prompting methods Wei et al. (2022); Wang et al. (2022); Zhou et al. (2022a), the explanations can be unfaithful Turpin et al. (2024); Lanham et al. (2023). Wiegreffe et al. (2022); Marasović et al. (2021) show potential for generating plausible free-text explanations with only a few examples, but the explanations can be unfaithful Ye and Durrett (2022); Chen et al. (2024).
Evaluate explanation faithfulness.
We do not assume access to ground-truth explanations for evaluation, reflecting real-world scenarios; therefore white-box faithfulness evaluation Zhou et al. (2022b); Chen et al. (2018) is not applicable to our problem. While robustness Alvarez-Melis and Jaakkola (2018) and perturbation-based evaluation methods Samek et al. (2016) have been applied in the vision domain, their applicability to NLP is limited due to the discrete nature of language inputs, and removing a single word can render a sentence meaningless Lyu et al. (2024). Simulatability-based methods have been employed to evaluate explanation Hase and Bansal (2020); Doshi-Velez and Kim (2017). This method is recommended to evaluate faithfulness Lyu et al. (2024), as the more faithful an explanation is, the more information it should contain about the model’s decision mechanism, and thus the easier it would be for an external simulator, to predict the model’s behavior based on the explanation. The assumption is that if an explanation leads to a different prediction than that made by the model it explains, then it is unfaithful Jacovi and Goldberg (2020). Pruthi et al. (2022) propose a framework for evaluating explanations by measuring accuracy gains in a student model trained to mimic a teacher. This framework is designed for system-level evaluation rather than individual-level, as the student model requires more than one examples to learn from.
3 Measuring and Improving Prediction-Explanation Consistency
Figure 2: Our framework for generating more consistent explanations. Given a question and a model predicted answer, we sample explanations from a language model and rank them using Prediction-EXplanation consistency (PEX), which measures how well an explanation $\bm{e}$ supports a given prediction $\bm{a}$ compared to its negation $\lnot\bm{a}$. This measure serves as a reward signal to construct preference dataset for direct preference optimization (DPO), improving PEX consistency. Finally, we evaluate whether the consistency-optimized explanations are more faithful.
Faithful explanations that accurately reflect a model’s reasoning Jacovi and Goldberg (2020); Lyu et al. (2024) promotes AI transparency and accountability. In this work, we aim to improve the faithfulness of free-text explanations by optimizing their consistency. To this end, we first introduce PEX (Prediction-EXplanation consistency) measure to quantify how well an explanation speaks in favor of (vs against) a prediction (§ 3.2). We then use this measure as an optimization criterion to generate more consistent explanations (§ 3.4), as illustrated in Figure 2. Finally, we evaluate the efficacy of this optimization in generating more faithful explanations (§ 4).
3.1 Problem: Inconsistent Explanations
Our first goal is to develop a measure of whether a model’s explanations are consistent in justifying its predictions $\bm{a}$ over an its negation $\lnot\bm{a}$, as a faithful explanation cannot simultaneously support both predictions Miller (2019); Lipton (1990). We observe that language model generated explanations can be inconsistent, e.g. in Figure 1, the explanation “use a lot of positive language” supports both the truthful and deceptive hypotheses, failing to distinguish why the model predicted answer $\bm{a}$ instead of the alternative prediction $\lnot\bm{a}$. As a result, the explanation is not consistent and, therefore, not faithful. We focus on generating explanations to justify opinion spam detection Ott et al. (2011); Hussain et al. (2020). We use this task because it requires minimal external knowledge while allowing for the generation of meaningful explanations.
We focus on language models that can answer questions, where a model $M$ takes a question $\bm{q}$ as input and generates an answer $\bm{a}$ as output with probability $M(\bm{a}\mid\bm{q})$. Additionally, we assume that upon request (e.g. prompting or other mechanisms) that the same model $M$ can generate a potentially faithful explanation of its prediction. We denote this as $M(\bm{e}\mid Q(\bm{q},\bm{a}))$, where $Q$ is an appropriate prompt, and explanation $\bm{e}=(e_{1},e_{2},...,e_{n})$ is a sequence of words. We use one-shot example to prompt the model to generate explanation for its prediction. The explanation usually contains 2-3 rationales. The explanations in the one-shot example are generated by GPT-4 (detailed in § A.1).
For example, in opinion spam detection, question $\bm{q}$ is formatted as: $\bm{q}=$ Is this review truthful or deceptive? Review: {review}. The output $\bm{a}$ is Truthful or Deceptive. If $M$ predicts the answer $\bm{a}$ as Truthful, the prompt $Q$ is formatted as: $Q(\bm{q},{\color[rgb]{0.35,0,0.35}\texttt{Truthful}})=$ Is this review truthful or deceptive? Review: {review}. Answer: Truthful. Question: Can you explain the main reasons why the review is truthful? If $M$ predicts the answer $\bm{a}$ as Deceptive, the prompt $Q$ is formatted as: $Q(\bm{q},{\color[rgb]{0.35,0,0.35}\texttt{Deceptive}})=$ Is this review truthful or deceptive? Review: {review}. Answer: Deceptive. Question: Can you explain the main reasons why the review is deceptive? An example of the generated explanation $\bm{e}$ is No specific examples: The review does not provide specific examples of the poor service or quality of the hotel.
3.2 Measuring Prediction-Explanation (PEX) Consistency
We define explanation inconsistency as failing to explicate why the model predicted $\bm{a}$ instead of its negation $\lnot\bm{a}$. To quantify this, we introduce Prediction-EXplanation consistency (PEX), which measures how well an explanation supports a given prediction compared to its negation. We build on the weight of evidence framework Melis et al. (2021); Wod (1985), extending it to to compute probabilities when evidence is given as sequences of words. The weight of evidence measures the extent to which an explanation supports or undermines a hypothesis, making it a reasonable measure of explanation consistency.
PEX consistency.
To measure the PEX consistency of explanation $\bm{e}$ for a given question $\bm{q}$ and model prediction $\bm{a}$, we compute a score to compare the likelihood of model $M$ generating $\bm{e}$ under different predictions:
$C(\bm{e})=\log\frac{M(\bm{e}\mid Q(\bm{q},\bm{a}))}{M(\bm{e}\mid Q(\bm{q},\lnot\bm{a}))}$ (1)
where $\lnot\bm{a}$ is a negation from $\bm{a}$, the text prompt $Q$ and explanation generation $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ are formatted in § 3.1. We estimate the conditional probability $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ using the chain rule:
$M(\bm{e}\mid Q(\bm{q},\bm{a})))=M(e_{1}\mid Q(\bm{q},\bm{a})))\cdot\\ M(e_{2}\mid Q(\bm{q},\bm{a})),e_{1})\cdot\\ \cdots M(e_{n}\mid Q(\bm{q},\bm{a})),e_{1},\ldots,e_{n-1})$ (2)
where $e_{i}$ is the $i$-th word of the explanation $\bm{e}$. We estimate the conditional probability $M(\bm{e}\mid Q(\bm{q},\lnot\bm{a}))$ using the same approach.
If we view $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ as $M(\bm{e}\mid\bm{q},\bm{a})$ by dropping the text format $Q$, PEX consistency $C(\bm{e})$ can also be computed using Baye’s rule as:
$\log\frac{M(\bm{e}\mid\bm{q},\bm{a})}{M(\bm{e}\mid\bm{q},\lnot\bm{a})}\\ =\log\frac{M(\bm{a}\mid\bm{q},\bm{e})}{M(\lnot\bm{a}\mid\bm{q},\bm{e})}-\log\frac{M(\bm{a}\mid\bm{q})}{M(\lnot\bm{a}\mid\bm{q})}$ (3)
where $\log\frac{M(\bm{a}\mid\bm{q},\bm{e})}{M(\lnot\bm{a}\mid\bm{q},\bm{e})}$ is the posterior log-odds ratio of between prediction $\bm{a}$ and $\lnot\bm{a}$ conditioned on question $\bm{q}$ and explanation $\bm{e}$. $\log\frac{M(\bm{a}\mid\bm{q})}{M(\lnot\bm{a}\mid\bm{q})}$ is the prior log-odds ratio, where $M(\bm{a}\mid\bm{q})$ is defined in § 3.1. To compute $M(\bm{a}\mid\bm{q},\bm{e})$ using a language model $M$, we format the prompt $Q^{\prime}$ as: $Q^{\prime}(\bm{q},\bm{e})$ = Is this review truthful or deceptive? Review: {review}. Analysis: $\bm{e}$. Then, we compute adjusted PEX consistency $C^{\prime}(\bm{e})$ as:
$C^{\prime}(\bm{e})=\log\frac{M(\bm{a}\mid Q^{\prime}(\bm{q},\bm{e}))}{M(\lnot\bm{a}\mid Q^{\prime}(\bm{q},\bm{e}))}-\log\frac{M(\bm{a}\mid\bm{q})}{M(\lnot\bm{a}\mid\bm{q})}$ (4)
We use adjusted PEX consistency (Eq 4) in our experiments because computing sequence probability for the original PEX consistency $C(\bm{e})$ (Eq 2), requires density estimation, which is often less reliable than computing classification probabilities for $C^{\prime}(\bm{e})$. Additionally, sequence probability is typically affected by sequence length.
Supervised fine-tuning.
We assume access to language models that provide output probabilities e.g. $M(\bm{a}\mid Q^{\prime}(\bm{q},\bm{e}))$, to compute PEX consistency (Eq 4). Pretrained language models, including LLama-2 and Mistra, do not perform well on opinion spam detection, with prediction accuracy close to random guessing. Thus we fine-tune the models on the training dataset using supervised fine-tuning training (SFT) to have better estimation of $M(\bm{a}\mid\bm{q})$. This is achieved by using the maximum-likelihood estimation (MLE) objective:
$\sum_{(q,a)\in D}\log M(\bm{a}\mid\bm{q})$ (5)
where the training dataset $D$ consisting of question answer pairs in the form of $(\bm{q},\bm{a})$. We use the fine-tuned models to measure PEX consistency.
3.3 Multiple Sampling for Generating More Consistent Explanations
The PEX measure (§ 3.2) can then be applied to rank language model generated explanations. For each question $\bm{q}$ and the answer $\bm{a}$ predicted by model $M$, we sample explanations $\bm{e}$ from $M$ using random sampling from the probability distribution $M(\bm{e}\mid Q(\bm{q},\bm{a}))$. We compute PEX consistency score (Eq 4) for each explanation, and rank the explanations according to the score.
3.4 Optimizing Explanation Consistency with Direct Preference Optimization
As the language models can generate inconsistent explanations for their predictions, we train models $M_{O}$ using the direct preference optimization (DPO) objective Rafailov et al. (2024) to generate more consistent explanations.
DPO training.
Given a question $\bm{q}$ and answer $\bm{a}$, $M_{O}$ is trained to increase the likelihood $M_{O}(\bm{e}_{w}\mid Q(\bm{q},\bm{a})))$ of generating consistent explanation $\bm{e}_{w}$ as preferred completion, and decrease the likelihood $M_{O}(\bm{e}_{l}\mid Q(\bm{q},\bm{a})))$ of generating inconsistent explanation $\bm{e}_{l}$ as dispreferred completion. We use the language model $M$ (§ 3.2) as base reference model, and use the DPO objective $\mathcal{L}(M_{O};M)$:
$\mathbb{E}_{(\bm{q},\bm{a},\bm{e}_{w},\bm{e}_{l})\sim D_{O}}\Bigg{[}\log\sigma\Big{(}\beta\log\frac{M_{O}(\bm{e}_{l}\mid Q(\bm{q},\bm{a}))}{M(\bm{e}_{l}\mid Q(\bm{q},\bm{a}))}\\ -\beta\log\frac{M_{O}(\bm{e}_{w}\mid Q(\bm{q},\bm{a}))}{M(\bm{e}_{w}\mid Q(\bm{q},\bm{a}))}\Big{)}\Bigg{]}$ (6)
where the weights of $M_{O}$ are initalized from the base reference model $M$. $\sigma$ is the logistic function, and $\beta$ is a parameter controlling the deviation from the base reference model $M$, set to 0.1.
Preference dataset.
To construct preference dataset $D_{O}$ for DPO training, we select $\bm{e}_{w}$ and $\bm{e}_{l}$ from the explanations generated by the reference language model $M$. For each question $\bm{q}$ and answer $\bm{a}$ predicted by model $M$, we sample and rank the explanations using PEX consistency score (Eq 4), as described in § 3.3. We consider explanations in the top $p\%$ are consistent, and those in the bottom $p\%$ are inconsistent. For each question and the answer, we use the combinations of the consistent and inconsistent explanations as preferred and dispreferred completions ($\bm{e}_{w}$, $\bm{e}_{l}$) to train DPO.
4 Explanation Faithfulness Evaluation
Figure 3: Explanation evaluation framework (figure reproduced from Pruthi et al., 2022): Student models learn from a teacher without explanation (top) or with explanation (bottom) as side information. Explanations $\bm{e}_{T}$ are effective if they improve student performance to mimic teacher model’s prediction on unseen examples during test phase without explanations (right).
To evaluate the efficacy of optimizing PEX consistency (§ 3) in generating more faithful explanations, we construct a proxy measure of faithfulness by assuming that a more faithful explanation should be more useful for an external observer to predict a model’s behavior (Lyu et al., 2024, see Section 3.4.2). Using this idea, we adopt a simulatability-based explanation faithfulness evaluation method Pruthi et al. (2022), as shown in Figure 3. In this framework, a teacher model generates a prediction and explanation, and a student model is trained to simulate the teacher’s model prediction. Explanations are available to the student model during training, but are not available during testing to avoid label leakage. The student model is models instead of humans, as human experiments are costly and humans’s preconceived notions may impair their ability to simulate the models accurately.
Student model training.
Our teacher model $M_{T}$ is a supervised fine-tuned model $M$ optimized for prediction accuracy (§ 3.2), and the student model $M_{S}$ is initialized from a pretrained language model. We fine-tune the student model $M_{S}$ to simulate the teacher’s model prediction and explanation: $M_{S}$ takes question $\bm{q}$ as input, and generates answer $\bm{a}$ and explanation $\bm{e}$ autoregressively as its response, formatted as: $R(\bm{a},\bm{e})={\color[rgb]{0.35,0,0.35}\texttt{Answer: $\bm{a}$. Analysis: $\bm{e}$}}$. The model is trained to estimate $M_{S}(R(\bm{a},\bm{e})\mid\bm{q})$ using the MLE objective:
$\sum_{(\bm{q},\bm{a}_{T},\bm{e}_{T})\in D_{T}}\>\sum_{n=1}^{N}logM_{S}(R_{n}\mid R_{<n},\bm{q})$ (7)
where $R_{n}$ is the $n$-th word of the text response $R(\bm{a},\bm{e})$, and $R_{<n}$ is the first $n-1$ words of the response. Dataset $D_{T}$ consisting of samples in the form of $(\bm{q},\bm{a}_{T},\bm{e}_{T})$, where $\bm{a}_{T}=M_{T}(\bm{q})$ is the answer predicted by the teacher model $M_{T}$, and explanation $\bm{e}_{T}$ is generated by $M_{T}$ to justify its prediction $\bm{a}_{T}$ to the question $\bm{q}$.
Evaluation.
During testing, the student model generates prediction $\bm{a}^{\prime}=M_{S}(\bm{q}^{\prime})$ for question $\bm{q}^{\prime}$, and we measure the simulation performance by comparing its prediction $\bm{a}^{\prime}$ with the teacher model’s prediction $\bm{a}_{T}^{\prime}=M_{T}(\bm{q}^{\prime})$.
5 Experimental Setup
Datasets.
We fine-tune pre-trained large language models to improve their prediction accuracy on two opinion spam detection datasets: (i) TripAdvisor hotel review dataset Ott et al. (2013), which contains 800 truthful reviews and 800 deceptive reviews. We randomly split the dataset to obtain 960 pairs of (review, label) for training, 320 pairs for validation and 320 pairs for testing. (ii) Amazon product review dataset Hussain et al. (2020), which we randomly select 1,000 truthful reviews and 1,000 deceptive reviews. We restrict reviews to those reviews containing at least 120 words to ensure that there is sufficient context for explanations. We split the selected reviews to obtain 1,200 pairs of (review, label) for training, 400 pairs for validation and 400 pairs for testing. We fine-tune the models on the train split, and select models on the validation split according to the F1 score.
We generate explanations using the fine-tuned models on the validation split to compute PEX consistency score statistics (§ 3.2). To evaluate explanation faithfulness, we also use the generated explanations on the validation split to serve as teacher model to fine-tune student model (§ 4). We report student model performance on the test split, evaluated against teacher predictions.
Pretrained-LLMs and SFT.
We evaluate three different LLM architectures to generate explanations for the opinion spam detection datasets: (i) Mistral-7B-Instruct-v0.3 Mistral (2023), (ii) Llama-2-13B-chat Touvron et al. (2023), and (iii) Yi-1.5-9B-chat Young et al. (2024). We fine-tune the pretrained models (§ 3.2), achieving an F1 score of 94% on the TripAdvisor validation set and 93% on the Amazon validation set. Details for fine-tuning the models are given in Appendix § A.2.
DPO training.
To construct the preference dataset for DPO training (§ 3.4), we generate 40 explanations per review-prediction pair in the training split (§ 3.3) using a sampling temperature of 1.0. To increase diversity, we sample 40 additional explanations at a temperature of 1.2. Explanations are ranked by their PEX consistency score (Eq 4), which measures how well they justify the prediction over its negation. For TripAdvisor dataset, we randomly sample 8 explanation pairs: preferred completions are drawn from the top 10% of scores and dispreferred from the bottom 10%, excluding pairs where both scores are above or below zero to ensure contrast. For Amazon dataset, we sample 8 pairs from the top and bottom 5% without applying a zero threshold, as many explanations for deceptive predictions have low scores. After this process, we obtain 3,565 training samples on the TripAdvisor dataset for DPO training with the Llama-2 model, 2,796 samples for the Mistral model, and 4,083 samples for the Yi-1.5 model. On the Amazon dataset, we obtain 9,600 training samples for each of the three language models. Additional details on model training are provided in § A.3.
Student model training.
We train student models  [^2] Our preliminary experiments do not suggest that using training-free in-context learning is fruitful. using $k$ random examples from the validation split, where each example includes both the teacher model’s prediction and its generated explanation (§ 4). We perform a five-pass training for $k=10$ and $k=20$. We use small values of $k$ because the student model can learn effectively from the prediction label alone when $k$ is large. The student model shares the same architecture as the teacher model but is not fine-tuned on the training split. Further details are provided in § A.2.
Evaluation metrics.
To measure how well the explanations justifies the prediction over its negation, we compute PEX consistency (Eq 4) on model generated explanations using the corresponding supervised fine-tuned model. To assess explanation faithfulness, we evaluate the trained student models using simulation performance (F1 score) on the test split, ensuring teacher model’s explanations are not provided as input to prevent label leakage (§ 4). The simulation F1 score is computed by using the teacher model’s predictions as ground-truth labels. We report average F1 score across all student model training passes.
6 Experiments
We investigate the following questions:
  1. 1.
    How consistent are the explanations generated by large language models (LLM)?
  2. 2.
    Can the consistency of LLM-generated explanations be improved?
  3. 3.
    Are explanations optimized for consistency also more faithful?
To address Q1, we analyze the PEX consistency score distribution of the explanations. For Q2, we train models using direct preference optimization to generate more consistent explanations. For Q3, we perform explanation faithfulness evaluation.
6.1 How consistent are the explanations generated by large language models?
Figure 4 shows the distribution of PEX consistency scores (§ 3.2) for explanations generated by the Mistral, Llama-2 and Yi-1.5 models on two opinion spam datasets. According to Wod (1985); Melis et al. (2021), weight of evidence scores above 2 are considered consistent. For the Mistral model, we see that 85.6% of explanations are inconsistent (with PEX score $<$ 0), 75.4% for the Llama-2 model, and 61.7% for the Yi-1.5 model. Since inconsistent explanations do not support the model’s predictions effectively, they pose a challenge for humans trying to understand how the model arrives at its decisions based on those explanations.
Figure 4: Prediction-EXplanation consistency (PEX) score distribution for different large language models. Explanations with PEX score higher than 2 are considered consistent Wod (1985). Takeaway: models can generate 62%-86% inconsistent explanations.
6.2 Can the consistency of LLM-generated explanations be improved?
(a) TripAdvisor dataset
(b) Amazon dataset
Figure 5: Explanation average PEX consistency scores for supervised fine-tuning (SFT) and direct preference optimization (DPO) models, computed on the TripAdvisor and Amazon validation datasets. The error bars represent 95% confidence intervals. DPO model scores are significantly higher than those of the SFT model ($p<0.05$ from two-related-sample t-test). Takeaway: DPO training can improve explanation consistency.
We use direct preference optimization (DPO) to enhance explanation consistency (§ 3.4), the results are illustrated in Figure 5. For the Mistral model, DPO improves consistency by 2.8 to 5.1 points compared to explanations generated from the supervised fine-tuning (SFT) model (§ 3.2). DPO also improves 2.6 to 6.6 points for the Llama-2 model, and 1.8 to 3.8 points for the Yi-1.5 model. These results demonstrate the effectiveness of DPO in improving the PEX consistency of the generated explanations, by learning the patterns that contribute to more consistent explanations. For the Mistral model on the TripAdvisor dataset, the average consistency scores of DPO-generated explanations remain below zero. This arises from the upper bound imposed by the sampled explanations being used to construct preference data: the highest consistency score among DPO-sampled explanations for this model averages -2.1 across the reviews, which is lower than the Llama-2 model’s highest score of 4.4. Similarly, the highest consistency score among DPO-sampled explanations for Mistral model averages -2.8 on the Amazon dataset, which is lower than the Yi-1.5 model’s highest score of 0.3.
6.3 Are explanations optimized for consistency also more faithful?
Model TripAdvisor Amazon
Mistral Llama-2 Yi-1.5 Mistral Llama-2 Yi-1.5
Pred Only 55.5 59.6 63.0 83.2 88.6 67.8
+ SFT 66.2 58.3 66.0 85.2 89.2 75.5
+ DPO 69.8 63.4 70.9 86.7 91.6 85.2
Table 1: Simulation performance (F1) of student models on the TripAdvisor and Amazon test set, evaluating how well they approximate the teacher model’s predictions on unseen examples. The student models are trained using explanations from different teacher models: supervised fine-tuning (SFT) or direct preference optimization (DPO) with PEX consistency measure. indicates results that are significantly higher than those of the SFT model, with $p<0.05$ as determined by a two-related-sample t-test. § A.4 shows confidence intervals across training passes. Takeaway: optimizing PEX consistency improves explanation faithfulness.
Figure 6: Direct preference optimization (DPO) generates better explanations: The Mistral SFT (supervised fine-tuning) model predicts the TripAdvisor review as truthful; however, its explanation includes the phrase “mentions an alternative hotel,” which is often associated with deceptive prediction, leading to inconsistency. In contrast, the DPO model produces more faithful explanations with a higher consistency score.
Systems.
We evaluate the faithfulness of explanations (§ 4) generated by two systems: (i) SFT, a supervised fine-tuning model (§ 3.2), and (ii) DPO, a direct preference optimization model that improves PEX consistency (§ 3.4). Additionally, we report simulation performance for Pred Only, which provides only the teacher model’s prediction without any explanations during student model training. We use the same review across all three systems to enable a fair comparison.
Consistency-optimized explanations improve explanation faithfulness.
Table 1 presents the student model’s simulation performance on the test set, where the student model was trained using explanations generated by different teacher models on the validation set. The F1 score measures how well the student model approximates the teacher model’s predictions on unseen data, averaged over different number of training samples (§ 5). On the TripAdvisor dataset, using DPO-generated explanations improves the F1 score by up to 14.3 points compared to not using explanations, and up to 5.1 points compared to using SFT explanations. The results indicate that explanations optimized using the PEX consistency measure with DPO better enable the student model to simulate the teacher’s predictions compared to explanations generated by the corresponding SFT model. On the Amazon dataset, using DPO-generated explanations improves the F1 score by up to 17.4 points compared to not using explanations, and up to 9.7 points compared to using SFT explanations. The DPO explanations from the Yi-1.5 model result in a larger F1 score gap compared to the SFT explanations, likely due to the model’s higher average PEX consistency score relative to the Mistral and Llama-2 models.
Qualitative examples.
Figure 6 presents an example where the explanation generated by the DPO model receives a higher PEX consistency score than the explanation generated by the SFT model. The SFT-generated explanation is also less faithful, as it often supports the deceptive prediction. Additional examples are provided in § A.5.
7 Conclusion
We present a new consistency measure for free-text explanations as an important aspect of faithfulness. We optimize explanation consistency using direct preference optimization, and show that it can improve the consistency and faithfulness of model generated explanations. We hope this research advances explanation faithfulness and enhances transparency in AI-assisted decision-making.
Limitations
Our approach to improving explanation consistency relies on computing the sequence output’s conditional probability from language models. Consequently, this approach cannot be directly applied to explanations generated by completely black-box models. We evaluate explanation faithfulness at the system level rather than at the individual explanation level, as the faithfulness evaluation framework we use is designed for system-level assessment. Our experiments focus on enhancing explanations for binary classification, but we envision future work extending this approach to multi-class classification, as weight of evidence can be applied in that scenario as well Melis et al. (2021).
Acknowledgements
This material is based upon work supported by the NSF under Grant No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), and U.S. Army Grant No. W911NF2120076. We thank Marine Carpuat, Navita Goyal, Amanda Liu, Connor Baumler, Ilaria Canavotto, John Horty, Eric Pacuit for discussion on explanation roles and evaluation. We also thank Nguyen X. Khanh for suggestion on direct preference optimization training, and Mingyang Xie for writing feedback.
Appendix A Appendices
Model Mistral Llama-2 Yi-1.5
$k$=10 $k$=20 Avg $k$=10 $k$=20 Avg $k$=10 $k$=20 Avg
Pred Only 44.0±7.8 66.9±1.5 55.5 54.7±0.1 64.2±1.4 59.6 62.4±0.0 63.6±0.0 63.0
+ SFT 64.2±1.8 68.1±0.2 66.2 56.0±4.0 60.6±2.2 58.3 66.5±0.0 65.4±0.7 66.0
+ DPO 69.1±1.2 70.4±1.0 69.8 61.4±0.9 65.3±2.9 63.4 69.3±0.1 72.5±0.1 70.9
Table 2: Simulation performance (F1) of student models on the TripAdvisor test set, evaluating how well they approximate the teacher model’s predictions on unseen examples. The student models are trained using explanations from different teacher models: supervised fine-tuning (SFT) or direct preference optimization (DPO) with PEX consistency measure. The variable $k$ represents the size of the teacher model’s training samples. The error bars after ± represent 95% confidence intervals across 5 training passes.
Model Mistral Llama-2 Yi-1.5
$k$=10 $k$=20 Avg $k$=10 $k$=20 Avg $k$=10 $k$=20 Avg
Pred Only 78.9±0.8 87.4±0.4 83.2 85.0±1.5 92.2±0.9 88.6 53.8±11.7 81.8±1.8 67.8
+ SFT 83.6±0.4 86.8±0.3 85.2 85.6±2.8 92.8±1.8 89.2 68.0±11.8 82.9±2.7 75.5
+ DPO 84.0±0.5 89.3±0.7 86.7 89.9±1.6 93.2±1.0 91.6 81.9±8.1 88.4±3.1 85.2
Table 3: Simulation performance (F1) of student models on the Amazon test set, evaluating how well they approximate the teacher model’s predictions on unseen examples. The student models are trained using explanations from different teacher models: supervised fine-tuning (SFT) or direct preference optimization (DPO) with PEX consistency measure. The variable $k$ represents the size of the teacher model’s training samples. The error bars after ± represent 95% confidence intervals across 5 training passes.
A.1 Explanation Prompting
We use the following GPT-4 generated explanation as one-shot example to prompt models for explanations (§ 3.1) for the TripAdvisor hotel review dataset:
Review: We stayed at Hotel Monaco and upgraded to a suite. I was disappointed with the room layout, and shabby furniture. I bumped into the coffee table and the top almost crashed to the ground. The make up mirror in the bathroom didnt work, the TV reception was very poor and overall the room just wasnt worth the \$450 we spent there. Next time, we will go back to Trump Tower across the street, which is far superior to Hotel Monaco. Explanation for truthful prediction: [reason1] Suggests an alternative hotel: The reviewer suggests an alternative hotel (Trump Tower) that they believe is superior, which is a practical and often truthful element in a hotel review. [reason2] Mentions specific details: The reviewer mentions specific details about the room, such as the room layout, shabby furniture. Review: My stay at the Hotel Monaco recently was fantastic. We got a very reasonable price for our stay (especially since we were a AAA member), and there was plenty of perks to utilize; in-room spa treatments and an evening wine wine hour to name a few. Another thing that the kids really enjoyed was that, upon request, the staff brought us a fish in its own fishbowl for our room. How cool is that! The staff were all very helpful and friendly and our stay was very enjoyable. I would definitely recommend the Hotel Monaco to anyone in the downtown Chicago area. Explanation for deceptive prediction: [reason1] Generic and overly positive tone: The review maintains a very positive tone without any specific details or occurrences that would substantiate such enthusiasm. The language used is quite general, lacking specific details or instances that typically accompany genuine experiences. [reason2] Uncommon or unrealistic details: The mention of receiving a fish in a fishbowl upon request is an uncommon and specific detail. While some hotels do offer unique services, the presentation of this detail, without further elaboration or context, seems exaggerated and implausible, casting doubt on the review’s authenticity.
We use the following GPT-4 generated explanation as one-shot example to prompt models for explanations (§ 3.1) for the Amazon product review dataset:
Review: I only wanted to use this to charge two phones from one adapter. This item at first seemed to be of decent quality. However, after 3 months of occasional use, the casing of one of the male plugs just fell apart in my hand as I was plugging it in. I did not apply undue force when it happened. After closer inspection, I could see how weak the plastic was -- tiny, brittle plastic clips hold the case together. I might have been able to glue or tape it back together, but given that it charges \$500 phones, not worth the risk of a short. There is a diode and and two wires soldered inside the housing that broke that could easilly short circuit. I do not recommend this product if you value the items you are charging. Explanation for truthful prediction: [reason1] Firsthand experience: The reviewer shares a personal account of using the product over three months, offering a credible, real-world perspective rather than speculation. [reason2] Safety concerns explained logically: The reviewer points out that the internal wiring and diode were exposed after the casing broke, increasing the risk of a short circuit. Review: I was looking to downsize from an otterbox defender (whose holster’s clip failed after one year) and this fits the bill nicely. Unlike the defender, the outside of the case is hard so it slides around (and in and out of your pocket) easily. You also don’t have to open and close littler rubber doors to gain access to the headphone jack and charging/syncing port. The only hassle is that the rubber lining around the headphone jack wouldn’t allow some of my older cheap cords or my good sony MDR 7506 headphones to plug in. Trimming the rubber solved the cheap cord problems but I had to replace the cord on the sonys with a hacked media bridge cord with a stepped plug on it. Explanation for deceptive prediction: [reason1] Minimizes major flaws: The review glosses over a serious issue — the case blocks headphone and charging ports — by treating it as a minor fix. [reason2] Shifts blame: Suggests the problem is due to ’older cheap cords’ rather than a design flaw, misleading readers.
A.2 Supervised Fine-tuning and Student Model Training
For supervised fine-tuning (SFT), we finetune Llama-2-13b-chat Touvron et al. (2023), Mistral-7B-Instruct-v0.3 Mistral (2023) and Yi-1.5-9B-chat Young et al. (2024) models to improve their prediction accuracy on two opinion spam classification datasets (§ 3.2).
Hyperparameters and computation.
For Llama-2, Mistral and Yi-1.5 models, we finetune our models for $20$ epochs, using a learning rate of $2e^{-4}$, batch size of $2$, and AdamW optimizer. We use a LoRA adapter Hu et al. (2021) of rank $64$ and alpha $16$. We finetune the models using $1$ RTXA6000 GPU, for $\sim 24$ hours. We train the student models for $100$ epochs using $10$ training examples and $50$ epochs using $20$ training examples, due to a small number of training samples. Training each model takes 1-2 hours.
Tools.
We implement our models with Pytorch 2.0.1, Huggingface Transformers 4.31.0, scikit-learn 1.2.2 and SciPy 1.15.1.
Datasets.
We two opinion spam detection datasets: (i) TripAdvisor hotel review dataset Ott et al. (2013) and (ii) Amazon product review dataset Hussain et al. (2020), under their Creative Commons Attribution-NonCommercial-ShareAlike 3.0 Unported License.
A.3 Direct Preference Optimization
Hyperparameters and computation.
For Llama-2, we finetune our models using a learning rate of $2e^{-4}$. For Mistral and Yi-1.5, we finetune our models using a learning rate of $2e^{-5}$. We finetune all the models for $5$ epochs on the TripAdvisor dataset, and $1$ epoch on the Amazon dataset. We use batch size of $2$, AdamW optimizer, a LoRA adapter Hu et al. (2021) of rank $32$ and alpha $16$. We finetune the models using $1$ RTXA6000 GPU, for $\sim 24$ hours.
Tools.
We implement DPO training using TRL 0.14.0.
A.4 Explanation Faithfulness Evaluation
Table 2 and Table 3 show the simulation performance for explanation faithfulness evaluation, broken down by different values of $k$, which represents the number of training samples being used to train the student model.
A.5 Additional Qualitative Examples
Figure 7 shows additional qualitative examples for the explanations generated by DPO and SFT.
Figure 7: DPO (Direct Preference Optimization) improves consistency score from the Llama-2 SFT (Supervised Fine-Tuning) model, but the generated explanation is still not consistent enough for its deceptive prediction on an Amazon product review.